OpenStudy (anonymous):

prove sin((a+b)/2) = cos (c/2)

OpenStudy (aravindg):

Well for a triangle a+b+c=180 so a+b=180-c sin((a+b)/2) here you can substitute a+b with 180-c sin(180-c))/2 =sin(90-c/2) =cos c/2

OpenStudy (anonymous):

ahhhh... i see i see.. thanks

OpenStudy (aravindg):

yw :)

OpenStudy (anonymous):

wait... why does the sin180 become sin90? :/

OpenStudy (aravindg):

(180-c))/2=180/2-c/2 =90-c/2

OpenStudy (aravindg):

gt it?

OpenStudy (anonymous):

... yeahhh...